OpenStudy (anonymous):

How do you evaluate the integral using the substitution method? 6.4e^(-3x+6)

hartnn (hartnn):

substitute u = -3x+6

OpenStudy (earthcitizen):

is that 6*4 or 6.40 ? \[\int\limits_{}^{} 64^{-3x+6} dx = -1/3\int\limits 64e^{u} * du = (64/3)*e^{-3x+6}\]

OpenStudy (earthcitizen):

...+C

OpenStudy (earthcitizen):

u get it, make u the subst. then differentiate it, chug it into dx and voila! vraiment!